OpenStudy (anonymous):

1. Amanda uses the equation y = kx to determine how much money (y) she earns after working any number of hours (x). Amanda earned $350 after working 40 hours. What is the value of k in Amanda’s equation? 2. At the North Fair you earn 9 points for each bull's eye you hit, but you lose 7 points for each miss. After 25 tries, Linda has 33 points. How many hits did she have?

OpenStudy (anonymous):

1. y = kx x= 40, y = 350 so 350 = k*40 , k = 35/4 = 8.75

OpenStudy (anonymous):

x+y = 25 x=hits y=misses 9x-7y=33 7x+7y = 175 add 16x = 208 x = 13 13 hits and 12 misses